30, Dude Ranch Circle estimates its uncollectible accounts at 1.5% of its credit sales of$825,000, when adjusting for estimated losses from uncollectible accounts, the debit to Bad Debt Expense is 31. The allowance for bad debts account is contra to which of the following accounts? ATan 32. After aging the accounts receivable, it is estimated that $1,200 will not be collected and the allowance account has an existing credit balance of $400. If the accounts receivable t net receivables would be otal $130,000, the 33. Under the allowance method, to write off an account that has been de the entry would jnclude determined to be unco llectible, to Mouon Ce eon 84. After aging the accounts receivable, it is estimated thąt $790 will hot the aging approach would be account has an existing debit balance of $230. The adjusting entry under for the amount of collected and the allowance the aging a t be nccaunts receiva ble, it is estimated th at $420 will not be collected and the allowance f $100. If accounts receivable is $145,000, the net receivables for Bad Debts
Answer #1

30. The debit of bad debt expense is $12375

Bad debt expense = credit sales * % of uncollectible = 825000*1.5% = 12375

31. Accounts receivable

It is accounts receivable to which the allowance for bad debts accounts is a contra to.

32. net receivables would be $1288001200 = 128800

33. debit to Allowance for Doubtful Accounts; credit to Accounts Receivable

Under the allowance method, the entry of uncollectible is written as follows:

Dr. Allowance for Doubtful Accounts

Cr. Accounts Receivable

34. the adjusting entry under the aging approach would be for the amount of $1020

(790+230) = 1020